### v3

# Uniformly $S$-pseudo-projective modules

Abstract.
In this paper, we introduce the notion of uniformly $S$-pseudo-projective ($u$-$S$-pseudo-projective) modules as a generalization of $u$-$S$-projective modules. Let $R$ be a ring and $S$ a multiplicative subset of $R$. An $R$-module $P$ is said to be $u$-$S$-pseudo-projective if for any submodule $K$ of $P$, there is $s\in S$ such that for any $u$-$S$-epimorphism $f:P\to\frac{P}{K}$, $sf$ can be lifted to an endomorphism $g:P\to P$. We prove that an $R$-module $M$ is $u$-$S$-quasi-projective if and only if $M\oplus M$ is $u$-$S$-pseudo-projective. Also, we prove that if $A\oplus B$ is $u$-$S$-pseudo-projective, then any $u$-$S$-epimorphism $f:A\to B$ $u$-$S$-splits. Several new characterizations of $u$-$S$-semisimple rings are given.
Key words and phrases:
$u$-$S$-projective, $u$-$S$-quasi-projective, $u$-$S$-pseudo-projective
2020 Mathematics Subject Classification:
13Cxx, 13C10, 13C12, 16D60.
(⋆) Corresponding author
1. Introduction
In this paper, all rings are commutative with a nonzero identity, and all modules are unitary unless otherwise stated. Let $R$ be a ring. A subset $S$ of $R$ is called a multiplicative subset of $R$ if $1\in S$, $0\notin S$, and $s_{1}s_{2}\in S$ for all $s_{1},s_{2}\in S$. Let $S$ be a multiplicative subset of a ring $R$ and $M,N,L$ be $R$-modules.
  1. (i)
    $M$ is called a $u$-$S$-torsion module if there exists $s\in S$ such that $sM=0$ [11].
  2. (ii)
    An $R$-homomorphism $f:M\to N$ is called a $u$-$S$-monomorphism ($u$-$S$-epimorphism) if $\text{Ker}(f)$ ($\text{Coker}(f)$) is a $u$-$S$-torsion module [11].
  3. (iii)
    An $R$-homomorphism $f:M\to N$ is called a $u$-$S$-isomorphism if $f$ is both a $u$-$S$-monomorphism and a $u$-$S$-epimorphism [11].
  4. (iv)
    An $R$-sequence $M\xrightarrow{f}N\xrightarrow{g}L$ is said to be $u$-$S$-exact if there exists $s\in S$ such that $s\text{Ker}(g)\subseteq\text{Im}(f)$ and $s\text{Im}(f)\subseteq\text{Ker}(g)$. A $u$-$S$ exact sequence $0\to M\to N\to L\to 0$ is called a short $u$-$S$-exact sequence [10].
  5. (v)
    A short $u$-$S$-exact sequence $0\to M\xrightarrow{f}N\xrightarrow{g}L\to 0$ is said to be $u$-$S$-split (with respect to $s$) if there is $s\in S$ and an $R$-homomorphism $f^{\prime}:N\to M$ such that $f^{\prime}f=s1_{M}$, where $1_{M}:M\to M$ is the identity map on $M$ [10].
W. Qi et al. [7] introduced the notion of $u$-$S$-injective modules. They defined an $R$-module $E$ to be $u$-$S$-injective if the induced sequence
$0\to\text{Hom}_{R}(C,E)\to\text{Hom}_{R}(B,E)\to\text{Hom}_{R}(A,E)\to 0$
is $u$-$S$-exact for any $u$-$S$-exact sequence $0\to A\to B\to C\to 0$. Dually, X. L. Zhang and W. Qi [10] introduced the notion of $u$-$S$-projective modules. They defined an $R$-module $P$ to be $u$-$S$-projective if the induced sequence
$0\to\text{Hom}_{R}(P,A)\to\text{Hom}_{R}(P,B)\to\text{Hom}_{R}(P,C)\to 0$
is $u$-$S$-exact for any $u$-$S$-exact sequence $0\to A\to B\to C\to 0$. They also introduced the notions of $u$-$S$-semisimple modules and $u$-$S$-semisimple rings. An $R$-module $M$ is called $u$-$S$-semisimple if any $u$-$S$-short exact sequence $0\to A\to M\to C\to 0$ is $u$-$S$-split. A ring $R$ is called $u$-$S$-semisimple if any free $R$-module is $u$-$S$-semisimple. Recently, M. Adarbeh and M. Saleh [2] introduced and studied the notion of $u$-$S$-injective ($u$-$S$-projective) relative to a module. They defined an $R$-module $M$ to be $u$-$S$-injective ($u$-$S$-projective) relative to a module $N$ if for any $u$-$S$-monomorphism $f:K\to N$ ($u$-$S$-epimorphism $g:N\to L$), the induced map $\text{Hom}_{R}(f,M):\text{Hom}_{R}(N,M)\to\text{Hom}_{R}(K,M)$ ($\text{Hom}_{R}(M,g):\text{Hom}_{R}(M,N)\to\text{Hom}_{R}(M,L)$) is a $u$-$S$-epimorphism. They also introduced the notion of $u$-$S$-quasi-injective ($u$-$S$-quasi-projective) modules. An $R$-module $M$ is called $u$-$S$-quasi-injective ($u$-$S$-quasi-projective) if it is $u$-$S$-injective ($u$-$S$-projective) relative to $M$. Afterwards, M. Adarbeh and M. Saleh [3] introduced and studied the notion of $u$-$S$-pseudo-injective modules. They defined an $R$-module $E$ to be $u$-$S$-pseudo-injective if for any submodule $K$ of $E$, there is $s\in S$ such that for any $u$-$S$-monomorphism $f:K\to E$, $sf$ can be extended to an endomorphism $g:E\to E$. In this paper, we introduce the notion of $u$-$S$-pseudo-projective modules as a dual notion of $u$-$S$-pseudo-injective modules. An $R$-module $P$ is said to be $u$-$S$-pseudo-projective if for any submodule $K$ of $P$, there is $s\in S$ such that for any $u$-$S$-epimorphism $f:P\to\frac{P}{K}$, $sf$ can be lifted to an endomorphism $g:P\to P$. From [2, Theorem 2.3], we conclude that an $R$-module $P$ is $u$-$S$-quasi-projective if and only if for any submodule $K$ of $P$, there is $s\in S$ such that for any $R$-homomorphism $f:P\to\frac{P}{K}$, $sf$ can be lifted to an endomorphism $g:P\to P$. From the last fact and [2, Remark 3.2 (1)], we have the following implications:
$u$-$S$-projective $\Rightarrow$ $u$-$S$-quasi-projective $\Rightarrow$ $u$-$S$-pseudo-projective.
In Section 2, we study some properties of $u$-$S$-pseudo-projective modules. Firstly, we show in Remark 2.2 (1) that the notions of $u$-$S$-pseudo-projective modules and pseudo-projective modules coincide under the condition that every element of $S$ is a unit of $R$. In Corollary 2.10, we show that an $R$-module $M$ is $u$-$S$-quasi-projective if and only if $M\oplus M$ is $u$-$S$-pseudo-projective. In Proposition 2.13, we show that if $M$ is an $R$-module and $f:P\to M$ is an epimorphism with $P$ projective, then $M$ is $u$-$S$-projective if and only if $P\oplus M$ is $u$-$S$-pseudo-projective. In Theorem 2.17, we show that a ring $R$ is strongly $S$-perfect if and only if every $u$-$S$-flat is quasi-projective if and only if every $u$-$S$-flat is pseudo-projective. Theorem 2.19 gives several new characterizations of $u$-$S$-semisimple rings. Theorem 2.21 characterizes rings in which every $u$-$S$-pseudo-projective module is $u$-$S$-projective.
Throughout, $U(R)$ denotes the set of all units of $R$; $\text{Max}(R)$ denotes the set of all maximal ideals of $R$; $\text{Spec}(R)$ denotes the set of all prime ideals of $R$.
2. $u$-$S$-pseudo-projective modules
Recall that an $R$-module $P$ is called pseudo-projective if for any submodule $K$ of $P$, every epimorphism $f:P\to\frac{P}{K}$ can be lifted to an endomorphism $g:P\to P$ [8]. We introduce the uniformly $S$-version of pseudo-projective modules.
Definition 2.1.
Let $S$ be a multiplicative subset of a ring $R$. An $R$-module $P$ is said to be $u$-$S$-pseudo-projective if for any submodule $K$ of $P$, there is $s\in S$ such that for any $u$-$S$-epimorphism $f:P\to\frac{P}{K}$, $sf$ can be lifted to an endomorphism $g:P\to P$.
Remark 2.2.
Let $S$ be a multiplicative subset of a ring $R$ and $P$ an $R$-module.
  1. (1)
    If $S\subseteq U(R)$, then $P$ is $u$-$S$-pseudo-projective if and only if $P$ is pseudo-projective.
  2. (2)
    $u$-$S$-projective $\Rightarrow$ $u$-$S$-quasi-projective $\Rightarrow$ $u$-$S$-pseudo-projective.
  3. (3)
    By (2) and [2, Proposition 3.6], every $u$-$S$-semisimple module is $u$-$S$-pseudo-projective.
For an $R$-module $M$, let $K\leq M$ denote that $K$ is a submodule of $M$. The following proposition gives some properties of $u$-$S$-pseudo-projective modules.
Proposition 2.3.
Let $S$ be a multiplicative subset of a ring $R$. Then the following statements hold:
  1. (1)
    Let $0\to A\xrightarrow{f}B\xrightarrow{g}C\to 0$ be a $u$-$S$-split $u$-$S$-exact sequence. If $B$ is $u$-$S$-pseudo-projective, then so are $A$ and $C$.
  2. (2)
    If $A\oplus B$ is $u$-$S$-pseudo-projective, then so are $A$ and $B$.
  3. (3)
    Let $f:A\to B$ be a $u$-$S$-isomorphism. Then $A$ is $u$-$S$-pseudo-projective if and only if $B$ is $u$-$S$-pseudo-projective.
  4. (4)
    If $A$ is a $u$-$S$-pseudo-projective module, then any $u$-$S$-epimorphism $f:A\to A$ $u$-$S$-splits.
Proof.
(1) Suppose that $0\to A\xrightarrow{f}B\xrightarrow{g}C\to 0$ is a $u$-$S$-split $u$-$S$-exact sequence. Then there are $R$-homomorphisms $f^{\prime}:B\to A$ and $g^{\prime}:C\to B$ such that $f^{\prime}f=t1_{A}$ and $gg^{\prime}=t1_{C}$ for some $t\in S$. Let $K\leq A$ and $H=f^{\prime-1}(K)$. Then $H\leq B$. Let $\eta_{K}:A\to\frac{A}{K}$ and $\eta_{H}:B\to\frac{B}{H}$ be the natural maps. Since $B$ is $u$-$S$-pseudo-projective, there is $s\in S$ such that for any $u$-$S$-epimorphism $h:B\to\frac{B}{H}$, there is $e\in\text{End}_{R}(B)$ such that $sh=\eta_{H}e$. Let $h^{\prime}:A\to\frac{A}{K}$ be any $u$-$S$-epimorphism. Define $\overline{f}:\frac{A}{K}\to\frac{B}{H}$ by $\overline{f}(a+K)=f(a)+H$. If $a-b\in K$, then $f^{\prime}(f(a)-f(b))=f^{\prime}f(a-b)=t(a-b)\in K$ and so $f(a)-f(b)\in f^{\prime-1}(K)=H$. Thus $\overline{f}$ is well-defined. It is easy to check that $\overline{f}$ is an $R$-homomorphism. Now if $b\in B$, then $f^{\prime}(b)\in A$ and so $tf^{\prime}(b)=f^{\prime}f(f^{\prime}(b))$. It follows that $f^{\prime}\big(tb-f(f^{\prime}(b))\big)=0\in K$. So $tb-f(f^{\prime}(b))\in f^{\prime-1}(K)=H$. Hence $t(b+H)=\overline{f}(f^{\prime}(b)+K)\in\text{Im}(\overline{f})$. Thus $\overline{f}$ is a $u$-$S$-epimorphism. Let $h:=\overline{f}h^{\prime}f^{\prime}:B\to\frac{B}{H}$. Then $h$ is a $u$-$S$-epimorphism since $f^{\prime},h^{\prime},\overline{f}$ are $u$-$S$-epimorphisms. So $sh=\eta_{H}e$ for some $e\in\text{End}_{R}(B)$. Let $e^{\prime}:=f^{\prime}ef$ and $s^{\prime}:=st^{2}$. Then $e^{\prime}\in\text{End}_{R}(A)$ and $st\overline{f}h^{\prime}=s\overline{f}h^{\prime}t1_{A}=s\overline{f}h^{\prime}f^{\prime}f=shf=\eta_{H}ef$. Let $a\in A$ and let $h^{\prime}(a)=a^{\prime}+K$. Then
$stf(a^{\prime})+H=st\overline{f}(a^{\prime}+K)=st\overline{f}h^{\prime}(a)=\eta_{H}ef(a)=ef(a)+H.$
So $stf(a^{\prime})-ef(a)\in H=f^{\prime-1}(K)$. It follows that $stf^{\prime}f(a^{\prime})-f^{\prime}ef(a)\in K$. So $st^{2}a^{\prime}-e^{\prime}(a)\in K$. Hence
$s^{\prime}h^{\prime}(a)=st^{2}a^{\prime}+K=e^{\prime}(a)+K=\eta_{K}e^{\prime}(a).$
Since $a$ was arbitrary, $s^{\prime}h^{\prime}=\eta_{K}e^{\prime}$. Thus $A$ is $u$-$S$-pseudo-projective. Similarly, we can show that $C$ is $u$-$S$-pseudo-projective.
(2) Let $i_{A}:A\to A\oplus B$ be the natural injection and $p_{B}:A\oplus B\to B$ be the natural projection. Since $0\to A\xrightarrow{i_{A}}A\oplus B\xrightarrow{p_{B}}B\to 0$ is a $u$-$S$-split $u$-$S$-exact sequence and $A\oplus B$ is $u$-$S$-pseudo-projective, then by part (1), $A$ and $B$ are $u$-$S$-pseudo-projective.
(3) This follows from part (1) and the fact that the $u$-$S$-exact sequences $0\to 0\to A\xrightarrow{f}B\to 0$ and $0\to A\xrightarrow{f}B\to 0\to 0$ are $u$-$S$-split.
(4) Suppose that $A$ is a $u$-$S$-pseudo-projective module. Let $f:A\to A$ be any $u$-$S$-epimorphism and let $K=\text{Ker}(f)$. Then $u:\frac{A}{K}\to A$ given by $u(a+K)=f(a)$, $a\in A$, is a $u$-$S$-isomorphism. By [10, Lemma 2.1], there is a $u$-$S$-isomorphism $v:A\to\frac{A}{K}$ and $t\in S$ such that $uv=t1_{A}$. Since $K\leq A$ and $A$ is $u$-$S$-pseudo-projective, then there is an endomorphism $e:A\to A$ such that $sv=\eta_{K}e$ for some $s\in S$. Thus $st1_{A}=suv=usv=u\eta_{K}e=fe$. Therefore, $f$ is $u$-$S$-split. ∎
Let $S$ be a multiplicative subset of a ring $R$ and $M$ an $R$-module. Recall that a submodule $N$ of $M$ is called a $u$-$S$-direct summand of $M$ if $M$ is $u$-$S$-isomorphic to $N\oplus N^{\prime}$ for some $R$-module $N^{\prime}$ [6].
Proposition 2.4.
Let $S$ be a multiplicative subset of a ring $R$. If $P$ is $u$-$S$-pseudo-projective and $K\leq P$ such that $\frac{P}{K}$ is isomorphic to a direct summand of $P$, then $K$ is a $u$-$S$-direct summand of $P$.
Proof.
Let $A$ be a direct summand of $P$ such that $\frac{P}{K}\cong A$. Let $f:A\to\frac{P}{K}$ be an isomorphism. If $p_{A}:P\to A$ is the projection map and $\eta_{K}:P\to\frac{P}{K}$ is the natural map, then $P\xrightarrow{p_{A}}A\xrightarrow{f}\frac{P}{K}$ is a $u$-$S$-epimorphism. Now since $P$ is $u$-$S$-pseudo-projective, there is $e\in\text{End}_{R}(P)$ such that the following diagram
[…]
commutes for some $s\in S$. So $sfp_{A}=\eta_{K}e$. Let $q:=ei_{A}f^{-1}:\frac{P}{K}\to P$, where $i_{A}:A\to P$ is the natural injection. Then
$\eta_{K}q=\eta_{K}ei_{A}f^{-1}=sfp_{A}i_{A}f^{-1}=sf1_{A}f^{-1}=sff^{-1}=s1_{\frac{P}{K}}.$
So by [10, Lemma 2.4], the exact sequence $0\to K\to P\to\frac{P}{K}\to 0$ is $u$-$S$-split. Hence by [6, Lemma 2.8], $P$ is $u$-$S$-isomorphic to $K\oplus\frac{P}{K}$. Therefore, $K$ is a $u$-$S$-direct summand of $P$. ∎
Corollary 2.5.
Let $S$ be a multiplicative subset of a ring $R$. Let $P$ be $u$-$S$-pseudo-projective and $e\in\text{End}_{R}(P)$. If $\text{Im}(e)$ is a direct summand of $P$, then $\text{Ker}(e)$ is a $u$-$S$-direct summand of $P$.
Proof.
Let $P$ be $u$-$S$-pseudo-projective and $e\in\text{End}_{R}(P)$. Then $\frac{P}{\text{Ker}(e)}\cong\text{Im}(e)$ is a direct summand of $P$. By Proposition 2.4, $\text{Ker}(e)$ is a $u$-$S$-direct summand of $P$. ∎
Let $S$ be a multiplicative subset of a ring $R$. Recall that $R$ is called $S$-von Neumann regular if for any $a\in R$, there exist $s\in S$ and $r\in R$ such that $sa=ra^{2}$ [11]. Equivalently, if the ring $R_{S}$ is von Neumann regular [11, Proposition 3.10]. Next, we extend the notion of $S$-von Neumann regular rings to noncommutative rings.
Definition 2.6.
Let $R$ be a noncommutative ring and $S$ a multiplicative subset of $R$. $R$ is said to be $S$-von Neumann regular if for any $a\in R$, there exist $s\in S$ and $r\in R$ such that $sa=ara$.
The following theorem gives sufficient conditions for the ring $\text{End}_{R}(P)$ to be $S$-von Neumann regular.
Theorem 2.7.
Let $S$ be a multiplicative subset of a ring $R$ and let $P$ be a $u$-$S$-pseudo-projective module such that $\text{Im}(e)$ is a direct summand of $P$ for every $e\in\text{End}_{R}(P)$. Then $\text{End}_{R}(P)$ is $S$-von Neumann regular ring.
Proof.
Let $e\in\text{End}_{R}(P)$. Then $\text{Im}(e)$ is a direct summand of $P$. By the proof of Proposition 2.4, the exact sequence $0\to\text{Ker}(e)\to P\xrightarrow{\eta}\frac{P}{\text{Ker}(e)}\to 0$ is $u$-$S$-split. So there is $s\in S$ and $R$-homomorphism $v:\frac{P}{\text{Ker}(e)}\to P$ such that $\eta v=s1_{\frac{P}{\text{Ker}(e)}}$. For $x\in P$, let $\overline{x}=x+\text{Ker}(e)$. Then $s\overline{x}=\eta v(\overline{x})=\overline{v(\overline{x})}$. So $sx-v(\overline{x})\in\text{Ker}(e)$, It follows that $sP\subseteq\text{Ker}(e)+\text{Im}(v)$. Also, we have $s\big(\text{Ker}(e)\cap\text{Im}(v)\big)=0$. Indeed, if $v(\overline{y})\in\text{Ker}(e)$, then $s\overline{y}=\eta v(\overline{y})=\overline{v(\overline{y})}=\overline{0}$ and so $sv(\overline{y})=0$. Let $P=A\oplus\text{Im}(e)$ be an internal direct sum. We have $se(P)\subseteq e(\text{Im}(v))$. Define $q:P\to P$ by $q(a+e(x))=sb$ where $a\in A$, $x\in P$, and $b\in\text{Im}(v)$ is such that $se(x)=e(b)$. Then $q$ is well defined. Indeed, if $a+e(x)=a^{\prime}+e(x^{\prime})$, then $a-a^{\prime}=e(x^{\prime}-x)\in A\cap\text{Im}(e)=0$. So $a=a^{\prime}$ and $x^{\prime}-x\in\text{Ker}(e)$. Write $se(x)=e(b)$ and $se(x^{\prime})=e(b^{\prime})$ for some $b,b^{\prime}\in\text{Im}(v)$. Since $e(b^{\prime}-b)=se(x^{\prime}-x)=0$, then $b^{\prime}-b\in\text{Ker}(e)\cap\text{Im}(v)$ and so $s(b^{\prime}-b)=0$. Hence $sb=sb^{\prime}$. Finally, we show that $s^{2}e=eqe$. Let $x\in P$. So $se(x)=e(b)$ for some $b\in\text{Im}(v)$. Thus $eqe(x)=e(q(0+e(x)))=e(sb)=se(b)=s^{2}e(x)$. Since $x\in P$ was an arbitrary, we have $s^{2}e=eqe$. Therefore, $\text{End}_{R}(P)$ is $S$-von Neumann regular ring. ∎
Lemma 2.8.
Let $S$ be a multiplicative subset of a ring $R$. Let $P$ be a $u$-$S$-pseudo-projective and $A$ a direct summand of $P$. Then for any $K\leq A$, there is $s\in S$ such that for any $u$-$S$-epimorphism $g:P\to\frac{A}{K}$, $sg$ can be lifted to a homomorphism $h:P\to A$.
Proof.
Let $P=A\oplus B$, $K\leq A$, and $H=K\oplus 0$. Identify $\frac{P}{H}=\frac{A\oplus B}{K\oplus 0}$ with $\frac{A}{K}\oplus B$. Since $P$ is $u$-$S$-pseudo-projective, there is $s\in S$ such that for any $u$-$S$-epimorphism $f:P\to\frac{P}{H}$, $sf$ can be lifted to an endomorphism $e:P\to P$. Let $g:P\to\frac{A}{K}$ be any $u$-$S$-epimorphism. Then the $R$-homomorphism $f:P\to\frac{P}{H}$ given by $f(a,b)=(g(a,b),b)$ is a $u$-$S$-epimorphism. So there is $e\in\text{End}_{R}(P)$ such that the following diagram
[…]
commutes. That is, $sf=\eta_{H}e$. Let $p_{A}:P\to A$, $p_{\frac{A}{K}}:\frac{A}{K}\oplus B\to\frac{A}{K}$ be the natural projections, and let $h:=p_{A}e:P\to A$. We claim that the following diagram
[…]
commutes. For $(a,b)\in P$, let $e(a,b)=(a^{\prime},b^{\prime})$. Then $\eta_{K}h(a,b)=\eta_{K}p_{A}e(a,b)=\eta_{K}(a^{\prime})=a^{\prime}+K=p_{\frac{A}{K}}(a^{\prime}+K,b^{\prime})=p_{\frac{A}{K}}\eta_{H}e(a,b)=p_{\frac{A}{K}}sf(a,b)=sg(a,b)$. Thus $sg=\eta_{K}h$. ∎
Proposition 2.9.
Let $S$ be a multiplicative subset of a ring $R$ and $A\oplus B$ be a $u$-$S$-pseudo-projective module. Then
  1. (1)
    AA is $u$-$S$-projective relative to $B$ and $B$ is $u$-$S$-projective relative to $A$.
  2. (2)
    any $u$-$S$-epimorphism $f:A\to B$ $u$-$S$-splits.
Proof.
(1) Suppose that $A\oplus B$ is $u$-$S$-pseudo-projective. Let $K\leq B$. Then by Lemma 2.8, there is $s\in S$ such that for any $u$-$S$-epimorphism $g:A\oplus B\to\frac{B}{K}$, $sg$ can be lifted to a homomorphism $h:A\oplus B\to B$. Let $f:A\to\frac{B}{K}$ be any homomorphism and $\eta_{K}:B\to\frac{B}{K}$ be the natural map. Define $g:A\oplus B\to\frac{B}{K}$ by $g(a,b)=f(a)+\eta_{K}(b)$. Then $g$ is an epimorphism and so it is a $u$-$S$-epimorphism. Hence $sg=\eta_{K}h$ for some homomorphism $h:A\oplus B\to B$. Let $i_{A}:A\to A\oplus B$ be the natural injection and $h^{\prime}:=hi_{A}:A\to B$. Then for $a\in A$,
$\eta_{K}h^{\prime}(a)=\eta_{K}hi_{A}(a)=sg(a,0)=s(f(a)+\eta_{K}(0))=sf(a).$
So $sf=\eta_{K}h^{\prime}$. Hence, the map $(\eta_{K})_{*}:\text{Hom}_{R}(A,B)\to\text{Hom}_{R}(A,\frac{B}{K})$ is a $u$-$S$-epimorphism. By [2, Theorem 2.3], $A$ is $u$-$S$-projective relative to $B$. Similarly, we can show that $B$ is $u$-$S$-projective relative to $A$.
(2) Let $f:A\to B$ be any $u$-$S$-epimorphism. Since $A\oplus B$ is $u$-$S$-pseudo-projective, then by part (1), $B$ is $u$-$S$-projective relative to $A$ and hence by [2, Lemma 2.15], the $u$-$S$-exact sequence $0\to\text{Ker}(f)\to A\xrightarrow{f}B\to 0$ $u$-$S$-splits. Thus $f$ $u$-$S$-splits. ∎
Corollary 2.10.
Let $S$ be a multiplicative subset of a ring $R$ and $M$ an $R$-module. Then $M$ is $u$-$S$-quasi-projective if and only if $M\oplus M$ is $u$-$S$-pseudo-projective.
Proof.
Suppose that $M$ is $u$-$S$-quasi-projective. Then $M\oplus M$ is $u$-$S$-quasi-projective by [2, Proposition 3.8 ]. So $M\oplus M$ is $u$-$S$-pseudo-projective by Remark 2.2 (2). The converse follows from Proposition 2.9 (1). ∎
Let $M$ be an $R$-module. For a positive integer $n$, let $M^{(n)}=\underbrace{M\oplus M\oplus\cdots\oplus M}_{n\text{-times}}$.
Corollary 2.11.
Let $S$ be a multiplicative subset of a ring $R$ and $M$ an $R$-module. For any integer $n\geq 2$, $M$ is $u$-$S$-quasi-projective if and only if $M^{(n)}$ is $u$-$S$-pseudo-projective.
Proof.
($\Rightarrow$). Since $M$ is $u$-$S$-quasi-projective, $M$ is $u$-$S$-projective relative to $M$. Hence by [2, Proposition 3.8], $M^{(n)}$ is $u$-$S$-quasi-projective and so by Remark 2.2 (2), $M^{(n)}$ is $u$-$S$-pseudo-projective.
($\Leftarrow$). For $n=2$, apply Corollary 2.10. For $n>2$, since $M^{(2)}\oplus M^{(n-2)}\cong M^{(n)}$ is $u$-$S$-pseudo-projective, then by Proposition 2.3 (2), $M^{(2)}$ is $u$-$S$-pseudo-projective. Thus $M$ is $u$-$S$-quasi-projective by Corollary 2.10. ∎
The following example gives a $u$-$S$-pseudo-projective module that is not $u$-$S$-projective.
Example 2.12.
Let $R=\mathbb{Z}$ and $S=\{1,2,3,\cdots\}$. Let $\mathbb{P}$ be the set of all prime numbers. Then $M:=\bigoplus\limits_{p\in\mathbb{P}}\mathbb{Z}_{p}$ is a $u$-$S$-quasi-projective module that is not $u$-$S$-projective by [2, Example 3.3]. By Corollary 2.10, $M\oplus M$ is a $u$-$S$-pseudo-projective module. However, $M\oplus M$ is not $u$-$S$-projective by [2, Corollary 2.8 (1)].
Proposition 2.13.
Let $S$ be a multiplicative subset of a ring $R$, $M$ an $R$-module, and $f:P\to M$ an epimorphism with $P$ projective. Then $M$ is $u$-$S$-projective if and only if $P\oplus M$ is $u$-$S$-pseudo-projective.
Proof.
Suppose that $M$ is $u$-$S$-projective. Since $P$ is projective, then by [10, Corollary 2.11], $P$ is $u$-$S$-projective. So by [10, Proposition 2.14 (1)], $P\oplus M$ is $u$-$S$-projective and hence by Remark 2.2 (2), $P\oplus M$ is $u$-$S$-pseudo-projective. Conversely, suppose that $P\oplus M$ is $u$-$S$-pseudo-projective. Then by Proposition 2.9 (2) and since $f:P\to M$ is a $u$-$S$-epimorphism, we have $f$ $u$-$S$-splits. So the exact sequence $0\to\text{Ker}(f)\to P\xrightarrow{f}M\to 0$ $u$-$S$-splits. Hence by [6, Lemma 2.8], $P$ is $u$-$S$-isomorphic to $\text{Ker}(f)\oplus M$. Since $P$ is $u$-$S$-projective, then by [10, Proposition 2.14 (3)] and [2, Corollary 2.8 (1)], $M$ is $u$-$S$-projective. ∎
Corollary 2.14.
Let $R$ be a ring, $M$ an $R$-module, and $f:P\to M$ an epimorphism with $P$ projective. Then $M$ is projective if and only if $P\oplus M$ is pseudo-projective.
For an $R$-module $M$, let $E(M)$ denote the injective envelope of $M$.
Proposition 2.15.
Let $S$ be a multiplicative subset of a ring $R$ and $M$ an $R$-module. Then $M$ is $u$-$S$-injective if and only if $M\oplus E(M)$ is $u$-$S$-pseudo-injective.
Proof.
Let $M$ be $u$-$S$-injective. Then $M\oplus E(M)$ is $u$-$S$-injective by [7, Proposition 4.7]. So by [3, Remark 2.4 (2)], $M\oplus E(M)$ is $u$-$S$-pseudo-injective. Conversely, suppose that $M\oplus E(M)$ is $u$-$S$-pseudo-injective. Then by [3, Theorem 2.6], the monomorphism $i_{M}:M\to E(M)$ $u$-$S$-splits. So the exact sequence $0\to M\to E(M)\to\frac{E(M)}{M}\to 0$ $u$-$S$-splits. Hence by [6, Lemma 2.8], $E(M)$ is $u$-$S$-isomorphic to $M\oplus\frac{E(M)}{M}$. But $E(M)$ is $u$-$S$-injective, so by [7, Proposition 4.7 (3)] and [2, Corollary 2.8 (2)], $M$ is $u$-$S$-injective. ∎
Corollary 2.16.
Let $R$ be a ring and $M$ an $R$-module. Then $M$ is injective if and only if $M\oplus E(M)$ is pseudo-injective.
Recall that an $R$-module $F$ is called $u$-$S$-flat if the induced sequence $0\to A\otimes_{R}F\to B\otimes_{R}F\to C\otimes_{R}F\to 0$ is $u$-$S$-exact for any $u$-$S$-exact sequence $0\to A\to B\to C\to 0$ [11], and that a ring $R$ is called a strongly $S$-perfect ring if every $u$-$S$-flat is projective [5].
Theorem 2.17.
Let $S$ be a multiplicative subset of a ring $R$. Then the following statements are equivalent:
  1. (1)
    $R$ is strongly $S$-perfect;
  2. (2)
    every $u$-$S$-flat is quasi-projective;
  3. (3)
    every $u$-$S$-flat is pseudo-projective.
Proof.
$(1)\Rightarrow(2)\Rightarrow(3)$. Clear.
$(3)\Rightarrow(1)$. Let $F$ be any $u$-$S$-flat $R$-module and $f:P\to F$ be an epimorphism with $P$ projective. Since $P$ is projective, it is $u$-$S$-flat by [10, Corollary 2.11 and Proposition 2.13]. So by [11, Proposition 3.4 (2)], $P\oplus F$ is $u$-$S$-flat. By (3), $P\oplus F$ is pseudo-projective, and so by Corollary 2.14, $F$ is projective. Hence, every $u$-$S$-flat $R$-module is projective. Therefore, $R$ is a strongly $S$-perfect ring. ∎
Theorem 2.18.
Let $S$ be a multiplicative subset of a ring $R$. Then the following statements are equivalent:
  1. (1)
    every finitely generated $u$-$S$-flat is projective;
  2. (2)
    every finitely generated $u$-$S$-flat is quasi-projective;
  3. (3)
    every finitely generated $u$-$S$-flat is pseudo-projective.
Proof.
Similar to the proof of Theorem 2.17. ∎
The following theorem gives several characterizations of $u$-$S$-semisimple rings.
Theorem 2.19.
Let $S$ be a multiplicative subset of a ring $R$. Then the following statements are equivalent:
  1. (1)
    $R$ is $u$-$S$-semisimple;
  2. (2)
    every $R$-module is $u$-$S$-pseudo-projective;
  3. (3)
    every direct sum of an $R$-module and a free $R$-module is $u$-$S$-pseudo-projective ($u$-$S$-pseudo-injective);
  4. (4)
    every direct sum of an $R$-module and a free $R$-module is $u$-$S$-quasi-projective ($u$-$S$-quasi-injective).
Proof.
$(1)\Rightarrow(2)$. Let $R$ be a $u$-$S$-semisimple ring. Then every $R$-module is $u$-$S$-quasi-projective by [2, Theorem 3.11]. So by Remark 2.2 (2), every $R$-module is $u$-$S$-pseudo-projective.
$(2)\Rightarrow(1)$. Let $M$ be any $R$-module and let $f:P\to M$ be an epimorphism with $P$ projective. Then by (2), $P\oplus M$ is $u$-$S$-pseudo-projective. So by Proposition 2.13, $M$ is $u$-$S$-projective. Hence every $R$-module is $u$-$S$-projective. Therefore, $R$ is $u$-$S$-semisimple by [10, Theorem 3.5].
$(1)\Rightarrow(3)$. This follows from $(1)\Rightarrow(2)$ and [3, Theorem 2.14].
$(3)\Rightarrow(1)$. Let $F$ be any free $R$-module. Then by (3), for any $R$-module $M$, $M\oplus F$ is $u$-$S$-pseudo-projective ($u$-$S$-pseudo-injective). By Proposition 2.9 (1) ([3, Proposition 2.13]), $M$ is $u$-$S$-projective ($u$-$S$-injective) relative to $F$ for any $R$-module $M$. Thus every $R$-module is $u$-$S$-projective ($u$-$S$-injective) relative to $F$, By [2, Theorem 2.16], $F$ is $u$-$S$-semisimple. Hence every free $R$-module is $u$-$S$-semisimple. Therefore, $R$ is $u$-$S$-semisimple.
$(1)\Leftrightarrow(4)$. This follows from [2, Theorem 2.16, Proposition 3.8, and Theorem 3.11]. ∎
Theorem 2.20.
Let $S$ be a multiplicative subset of a ring $R$. Then the following statements are equivalent:
  1. (1)
    Every $u$-$S$-quasi-projective module is $u$-$S$-projective;
  2. (2)
    Every direct sum of two $u$-$S$-quasi-projective modules is $u$-$S$-projective;
  3. (3)
    Every direct sum of two $u$-$S$-quasi-projective modules is $u$-$S$-quasi-projective.
Proof.
$(1)\Rightarrow(2)$. Let $M$ and $N$ be two $u$-$S$-quasi-projective modules. Then by (1), $M$ and $N$ are $u$-$S$-projective modules. So $M\oplus N$ is $u$-$S$-projective by [10, Proposition 2.14 (1)].
$(2)\Rightarrow(3)$. This is clear from Remark 2.2 (2).
$(3)\Rightarrow(1)$. Let $M$ be a $u$-$S$-quasi-projective module. Let $f:P\to M$ be an epimorphism with $P$ projective. Then by (3), $P\oplus M$ is $u$-$S$-quasi-projective. Let $p_{1}:P\oplus M\to P$ and $p_{2}:P\oplus M\to M$ be the natural projections, and let $i_{2}:M\to P\oplus M$ be the natural injection. Since $P\oplus M$ is $u$-$S$-quasi-projective and $fp_{1}:P\oplus M\to M$ is an epimorphism, then by [2, Theorem 2.3], there is $e\in\text{End}_{R}(P\oplus M)$ such that the following diagram
${P\oplus M\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{sp_{2}}$${P\oplus M\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{e}$$\scriptstyle{fp_{1}}$${M}$
commutes for some $s\in S$. So $sp_{2}=fp_{1}e$. Since $p_{2}i_{2}=1_{M}$, then $s1_{M}=sp_{2}i_{2}=fp_{1}ei_{2}$. Let $q:=p_{1}ei_{2}:M\to P$. Then $s1_{M}=fq$. So by [10, Lemma 2.4], the exact sequence $0\to\text{Ker}(f)\to P\xrightarrow{f}M\to 0$ $u$-$S$-splits. By [6, Lemma 2.8], $P$ is $u$-$S$-isomorphic to $\text{Ker}(f)\oplus M$. Since $P$ is $u$-$S$-projective, $\text{Ker}(f)\oplus M$ is $u$-$S$-projective by [10, Proposition 2.14 (3)]. Hence $M$ is $u$-$S$-projective by [2, Corollary 2.8 (1)]. Thus (1) holds. ∎
The following result gives a characterization of rings in which every $u$-$S$-pseudo-projective module is $u$-$S$-projective.
Theorem 2.21.
Let $S$ be a multiplicative subset of a ring $R$. Then the following statements are equivalent:
  1. (1)
    Every $u$-$S$-pseudo-projective module is $u$-$S$-projective.
  2. (2)
    Every direct sum of two $u$-$S$-pseudo-projective modules is $u$-$S$-projective.
  3. (3)
    Every direct sum of two $u$-$S$-pseudo-projective modules is $u$-$S$-pseudo-projective.
Proof.
$(1)\Rightarrow(2)$. Let $M$ and $N$ be two $u$-$S$-pseudo-projective modules. Then $M$ and $N$ are $u$-$S$-projective modules and so $M\oplus N$ is $u$-$S$-projective.
$(2)\Rightarrow(3)$. This follows from Remark 2.2 (2).
$(3)\Rightarrow(1)$. Let $M$ be any $u$-$S$-pseudo-projective module. Let $f:P\to M$ be an epimorphism with $P$ projective. Then by (3), $P\oplus M$ is $u$-$S$-pseudo-projective. Thus $M$ is $u$-$S$-projective by Proposition 2.13. ∎
Conflict of interest: The authors declare that they have no conflict of interest.